if you stand on a pogo stick you note that its spring loaded foot is pushed in about 18cm. if you bounce gently up and down, what is your approximate oscillation frequency? The answer is 1.2 Hz but please show work
Answer #1

as we know

mg = kx

g/x = k/m

as we know

w^2 = k/m

so

w^2 = g/x

(2pi f)^2 = 9.8 / 0.18

f = 7.379 / (2* 3.14)

f = 1.175 Hz

f = 1.2 Hz

=====

Comment before rate in case any doubt, will reply for sure.. goodluck